### v1

# Affordable Generative Agents

Abstract
The emergence of large language models (LLMs) has significantly advanced the simulation of believable interactive agents. However, the substantial cost on maintaining the prolonged agent interactions poses challenge over the deployment of believable LLM-based agents. Therefore, in this paper, we develop Affordable Generative Agents (AGA), a framework for enabling the generation of believable and low-cost interactions on both agent-environment and inter-agents levels. Specifically, for agent-environment interactions, we substitute repetitive LLM inferences with learned policies; while for inter-agent interactions, we model the social relationships between agents and compress auxiliary dialogue information. Extensive experiments on multiple environments show the effectiveness and efficiency of our proposed framework. Also, we delve into the mechanisms of emergent believable behaviors lying in LLM agents, demonstrating that agents can only generate finite behaviors in fixed environments, based upon which, we understand ways to facilitate emergent interaction behaviors. Our code is publicly available at: https://github.com/AffordableGenerativeAgents/Affordable-Generative-Agents.
Machine Learning, ICML
1 Introduction
Constructing an agent based on large language models (LLMs) for the simulation of believable human behavior holds substantial promise (da Rocha Costa, 2019; Park et al., 2022; Xi et al., 2023a). A believable LLM agent is characterized by its ability to interact with other agents or humans, respond to environmental changes, and generate responses and behaviors that are perceived by humans as authentic, natural, and in alignment with expectations. Such agents could be utilized for diverse applications, such as simulating the potential impacts of policies (Zhang et al., 2023a; Törnberg et al., 2023) or sociological theories on humans (Gao et al., 2023), creating Non-Player Characters (NPCs) (Laird & VanLent, 2001; Gong et al., 2023) in games, and developing embodied intelligence akin to social robots (Cherakara et al., 2023) or domestic service robots (Puig et al., 2020; Yang et al., 2023b).
However, while we appreciate the potential and capabilities of these agents, their cost implications also merit attention. The crux of the matter is that real-time, long-term invocation of LLMs to handle various types of agent interactions can result in substantial cost consumption. Previous studies have highlighted this issue. For instance, Park et al. (2023) construct a virtual town, aka the “Stanford Town”, to simulate social behavior, with the entire study incurring costs of thousands of dollars. Wang et al. (2023b) mimics embodied agents, fulfilling both emotional and rational needs through interaction, with simulations of 2 to 3 agents costing several dollars per hour.
The cost of believable agent interaction has hindered its use as a low-cost solution for simulating policy impacts and sociological phenomena, especially in long-term, large-scale scenarios. The cost, which increases linearly with the number of interactions, is unacceptable, particularly when LLM agents are to be used as NPCs in an open world environment.
To address the issues above, there has been a significant amount of research (Liang et al., 2020; So et al., 2021; Miao et al., 2023; Chen et al., 2023a; Zhang et al., 2023b). Nevertheless, existing methods primarily focus on specific independent tasks, solving problems through single-round QA sessions with LLMs (Chen et al., 2023a; Zhang et al., 2023b). Consequently, there lacks solutions for continuous multi-round agent interactions and the coordination infrastructure supporting low-cost implementations. To expand, the main reasons these methods are unsuitable for believable behavior simulation include the following aspects: 1) the lack of a unified and explicit evaluation system for believable behavior in open-world scenarios, making it unsuitable for task difficulty segmentation and cost-efficient model inference (Khalili et al., 2022); 2) the need for agents to consider context and multi-round interactions, requiring large amounts of relevant information for decision-making, making it unsuitable for combining multiple independent tasks into a single prompt for single-round inference (Lin et al., 2023; Mu et al., 2023); and 3) the need to consider the cost of single interactions as well as the redundancy in large-scale interactions, which requires optimization from the framework level.
Therefore, in this paper, we focus on the believable behavior interaction simulation of LLM agents in open-world scenarios and propose a low-cost framework, termed as Affordable Generative Agents. For the agent-environment interactions, we propose the Lifestyle Policy to reduce the redundant costs of generating repeated responses by agents; On the other hand, for the inter-agent interactions, we propose the Social Memory to reduce the generation of repeated information in multi-round interactions. These two techniques are distilled from our insights on how a human interacts with other entities in the real world, and can be compiled easily into existing agent simulation environments. To examine the applicability of our techniques, we have conducted extensive experiments using well-known environments, including the “Stanford Town” (Park et al., 2023) and the VirtualHome (Puig et al., 2018), to demonstrate that, while achieving the same performance, the consumption of generating believable agent behaviors can be significantly reduced.
Furthermore, to help understand why our approach works, we delve into the mechanics of how LLM agents generate believable behavior, demonstrating that there is an upper limit to the believable behavior that emerges from LLM agents in a fixed environment. This implies that all agent-environment interactions can be entirely covered by policies. Based upon this finding, we also propose a set of optimization strategies for generating richer believable behavior.
In summary, our contributions are as follows:
  • We propose Affordable Generative Agents, a low-cost simulation framework for the believable behavior simulation of LLM agents in open-world scenarios, with optimization strategies for both agent-environment and inter-agent interactions.
  • We propose several evaluation methods and conduct extensive experiments in benchmarking environments to validate the effectiveness of our framework.
  • We analyze the mechanism of believable behavior generation by LLM agents, demonstrate the upper limit of the richness of emergent behaviors, and further propose corresponding optimization strategies.
2 Related Work
This work explores the development of efficient generative behaviors leveraging LLMs. We discuss the most related works below.
2.1 LLM Agents
The development of LLM agents has sparked a myriad of applications. Yang et al. (2023a) leverages multi-turn dialogues with agents to autonomously address NLP tasks. Deng et al. (2023), Nakano et al. (2021), and Yao et al. (2022a) integrate pre-defined APIs with LLMs to enable agents to navigate web pages effectively. Xu et al. (2023) and Light et al. (2023) deploy LLM agents in text game environments, showing the capacity to engage in complex, strategic gameplay. The most prominent among these works is the construction of LLM agents that interact in open environments to create believable behavior. For instance, Park et al. (2023) builds a virtual town simulating the social life of 25 agents, while Brohan et al. (2023) utilizes LLM to create a robot capable of completing household chores through natural language interaction. Wang et al. (2023a) implements an agent capable of playing Minecraft expertly. Besides, researchers have been focusing on enhancing its performance. Efforts have been made to optimize various modules such as reasoning (Wei et al., 2022; Wang et al., 2022; Xi et al., 2023b), planning (Song et al., 2023; Yao et al., 2022b), reflection (Shinn et al., 2023), and memory (Zhong et al., 2023; Huang et al., 2023b), or to create more realistic interaction environments (Hong et al., 2023b; Huang et al., 2023a). However, the development of affordable and believable agent interactions is missing in the literature. In this paper, we focus on achieving believable agents with the same performance at low cost.
2.2 The Cost of LLM Interaction
The cost of using LLM is directly proportional to the number of LLM invocations, the length of the prompt, and the model’s performance. The cost issue has always been a key focus in the LLM-related field. Numerous methods have been proposed to reduce the cost of using LLM, such as Cascaded Invocation of LLM (Yue et al., 2023; Chen et al., 2023a; Nottingham et al., 2023), summary (Laban et al., 2023; Arefeen et al., 2023; Mu et al., 2023), and batch (Lin et al., 2023). These methods primarily optimize the cost of single invocation of LLM for QA tasks. Zhang et al. (2023b) has built a cost-efficient multi-agent framework, but it still solves QA tasks on fixed datasets. The simulation of believable behavior in the open world requires uninterrupted calls to the LLM for interaction. Moreover, multiple modules need to combine a large amount of contextual information for LLM reasoning and planning, making cost a more prominent issue. Simply optimizing for single QA invocations cannot effectively reduce costs. Furthermore, the open world does not have a fixed evaluation method, and the difficulty of tasks is hard to categorize, rendering many existing methods unsuitable (Chen et al., 2023a; Zhang et al., 2023b). Kaiya et al. (2023) proposes generative agents with low-cost interactions, but their effectiveness is highly dependent on the tailored scenarios. Hence, there is a pressing need for a low-cost solution that can effectively handle the challenges of believable interaction simulation in various open environments using LLMs. This paper aims to address this gap by proposing a cost-efficient scheme for simulating believable behavior with LLMs in open environments.
3 Method
Figure 1: The AGA framework, including (a) Agent-environment interactions and (b) Inter-agent interactions.
To implement the AGA framework, we abstract the interaction patterns of the LLM agent into two types: agent-environment interactions and inter-agent interactions. Figure 1 provides an overview of the AGA framework. In the following sub-sections, we will describe the optimization strategies for these two categories of interactions in detail.
3.1 Agent-environment interactions
The interaction between an agent and its environment can be modeled as a policy function mapping observations to actions. This policy can be implemented through reinforcement learning (Kaelbling et al., 1996) or rule-based finite state machines (Lee & Yannakakis, 1996), behavior trees (Colledanchise & Ögren, 2018), etc. LLMs trained on a vast amount of data show immense potential in policy formulation (Sun et al., 2023). With carefully designed prompts, LLMs can generate logically consistent and believable decisions. The decision-making capabilities of LLMs have been applied to a wide variety of applications, but the cost associated with each interaction has been overlooked. On the other hand, behavior trees operate without additional costs, but they require pre-written scripts and can not adapt dynamically to environment changes.
To combine the advantages of these two methods, we propose a new method – Lifestyle Policy, which aims to retain the flexible interaction capabilities of the LLMs while achieving low-cost deployment. The functionality of Lifestyle Policy is to minimize costs by reusing the similar inference processes of LLM. To this end, we devise agent-environment interactions containing two stages: Plan Decomposition and Policy Reuse, expanded as follows.
Plan Decomposition
The stage occurs when the agent encounters situations it has never seen before. Due to their inherent zero-shot capabilities (Kojima et al., 2022), LLM agents can respond to unseen tasks in a believable way. Given the description of the environment and agent’s information, a rough plan or task is generated by LLM. The Plan Decomposition breaks the plan down into sub-plans and then converts sub-plans into specific actions, just like other LLM multi-agent framework (Chen et al., 2023b; Li et al., 2023; Hong et al., 2023a). An additional step involves generating an executable condition based on the action sequence. The determination of execution condition is highly correlated with the environment. For instance, in the case of a domestic robot (Brohan et al., 2023), the condition of a task include the existence of the related interactive items and the consistency of their states. If the plan can be successfully executed, then the whole graph from plan to actions and corresponding condition will be updated to Lifestyle Policy.
Policy Reuse
The stage occurs when the agent encounters previously seen situations. Given a plan, we first search the Lifestyle Policy for any existing plans that match or closely resemble the new one, based on cosine similarity of their embedding features. When the similarity between two plans exceeds a certain threshold, we consider them to have same meaning, albeit with minor differences in their descriptions. Subsequently, we assess whether the current environment meets the execution condition associated with the retrieved plan. If the condition is met, we proceed with the actions specified in that plan. It is important to note that a single plan may have multiple conditions tailored for different scenarios. We iterate through all the conditions until we find one that the current environment satisfies.
Figure 1(a) illustrates the complete interaction process. Upon receiving an observation that includes environmental data and the agent’s personal profile, the LLM agent formulates a plan. Initially, the Lifestyle Policy retrieves a plan with a similar description and condition met by current environment. If a match is found, the agent executes the corresponding actions. Otherwise, the LLM is invoked to decompose the plan. The LLM plan decomposition ensures that the actions generated are of high quality, believable, and self-consistent. Meanwhile, the Lifestyle Policy aims to reuse existing plans to avoid unnecessary costs.
In the lifelong simulation of LLM agents, there are a large number of redundant, repetitive, and trivial decisions, such as in Generative agents (Park et al., 2023), where each agent repeatedly makes decisions like brushing teeth, washing face, etc. Replacing the process of invoking LLM reasoning with corresponding policies can save lots of cost, which will be shown in our experiments.
3.2 Inter-agent interactions
In agent interactions, dialogue is the primary communication method. These agents must integrate substantial auxiliary information into the dialogue prompt to demonstrate advanced cognitive capabilities, distinguishing them from basic chatbots driven by LLMs. For instance, agents need to incorporate relevant information about the dialogue participants and topic-related events retrieved from memory into the dialogue prompt to enrich the conversation. However, this data is often less comprehensive and more fragmented than the knowledge bases of question-answering systems (Cui et al., 2023). Merely appending memory snippets to prompts could reduce response quality and lead to higher processing costs from increased prompt length (Krishna et al., 2022). Furthermore, agent interactions differ depending on their relationships and mutual perceptions in social simulations. Thus, agents need to gather enough information to form contextually appropriate responses without suffering performance drops or incurring extra costs from excessively long prompts.
Drawing inspiration from existing text compression technologies (Liu et al., 2023), we introduce the Social Memory module, aimed at reducing the cost of inter-agent interactions and enabling the agents to produce more appropriate dialogues. This module is comprised of three key components: Relationship, Feeling, and Summary Events. Together, these components model the characteristics involved in agent interactions. The relationship and feeling components are responsible for recording the social connections and personal impressions between agents. These records shape how the agent interacts with others, influencing their behavior and approach. they can be a word or a phrase, like “Relationship: acquaintances, Feeling: friendly”. Notably, the two components are updated after each interaction to reflect any potential changes in relationships and emotions. The summary events component, on the other hand, focuses on summarizing all events related to the interlocutors and the topics of discussion. These events are retrieved from agent’s memory, enabling agents to quickly access necessary information from condensed text, thereby enhancing interaction efficiency.
Figure 1(b) presents the procedure of a dialogue interaction. Given the previous round of conversation, the agent computes the dialogue’s embedding features and retrieves relevant events from memory. After filtering out duplicates, these events are consolidated into summary events. As more interactions occur, additional relevant events will be integrated into the summary events module, which aims to compress information densely within a constrained text area, improving the efficiency and effectiveness of information delivery. The dialogue prompt is then formed by merging information from the last round of conversation and Social Memory, and the LLM is invoked generates the next round of dialogue. Upon the conclusion of the conversation, the relationship and feeling are updated based on the content of the dialogue.
Social Memory reduces costs of conversation by compressing auxiliary information, offering precise descriptions of social information to assist agents in responding consistently and appropriately. It also supplies high-information events to aid agents in generating accurate replies relevant to the conversation. Furthermore, as illustrated in Figure 2, Social Memory updates after each interaction, providing an explicit depiction of interlocutor relationships. This mechanism can be employed to investigate the dynamics of relationships between agents, thereby analyzing the evolution of community relations and the emergence of social behaviors.
Figure 2: An example of relational evolution driven by the updating of Social Memory following conversational interactions.
4 Experimental Setup
This section focuses on the experimental environment and evaluation protocol.
4.1 Environment
We evaluate our approach under the following environments:
  • Generative Agents. Park et al. (2023) creates a simulated town, populating it with multiple agents powered by LLMs that interact in natural language, generating believable individual and emergent social behaviors. Agents are required to create daily schedules, which are refined into hourly plans, and then translated into minute-level activities. All interactions are executed through natural language descriptions, accompanied by corresponding emojis for representation.
    The Generative Agents framework comprises five modules: Perceive, Plan, Retrieve, Reflect, and Act. The Perceive module is responsible for sensing surrounding game objects. The Plan module generates plans and activities. Retrieve is tasked with retrieving relevant events from memory. Reflect deeply considers past events, and Act determines the location and objects of interaction, as well as generating corresponding emojis.
    We implement AGA on Generative Agents framework and conduct experiments in the provided 3-person and 25-person environments.
  • VirtualHome. Puig et al. (2018) provides a 3D simulated domestic setting where agents engage with their surroundings via programmatic atomic actions. The environment encompasses 115 specific types of items, such as plate, desk, and microwave. Each item possesses attributes that determine the permitted interactions, including GRABBABLE, LIEABLE, etc., as well as current states like OPEN and CLOSED. Compared to Generative Agents, VirtualHome offers more concrete and realistic interaction. An interaction involves specific actions, items and the item states.
    We implement a LLM agent framework like Generative Agents on VirtualHome to validate our method.
4.2 Language Models Adopted
For a fair comparison, all agents are empowered with GPT-3.5-Turbo (Wu et al., 2023), same as the Generative Agents work (Park et al., 2023). In the evaluation phase, we use GPT-4 (Achiam et al., 2023) as an evaluator to assess the agent’s believable behavior.
4.3 Evaluation Protocol
We focus on token consumption and performance. We carry out a series of experiments to calculate the mean and standard deviation of token consumption. To comprehensively evaluate the agent’s behavior, we first reuse the evaluation methods of Generative Agents excluding manual assessments. These methods encompass interviewing agents with manually designed questions and conducting case study-specific activities.
Moreover, we propose two novel methods to evaluate the emergent social behaviors of generative agents:
  • Relationship Evolution. The relationships among agents are updated by Social Memory after interactions. The evolution of these relationships signifies the formation of community ties, which can be utilized to examine the social behaviors.
  • LLM Evaluator. Chiang & Lee (2023) have validated LLM as an evaluator, demonstrating comparable performance to expert evaluators. Consequently, we employ GPT-4 to evaluate the activities and dialogues generated by LLM agents. A 5-point Likert scale questionnaire is deployed to differentiate whether these activities and dialogues are from human or AI-generated.
For VirtualHome, agents are designed to experience a full day at home, generating household tasks and decomposing them into specific actions. We evaluate the success rate of plan execution.
5 Evaluations
We conduct extensive experiments to validate our method, assessing both the token consumption and the performance. Our experiments demonstrate that the AGA framework can reduce the cost while maintaining equivalent performance.
5.1 Results on Generative Agents
Token Consumption
Figure 3 illustrates the ablation study of token consumption based on multiple two game day simulations of 3-person town. Compared to the baseline, the cost of using only the Lifestyle Policy is 40.2% of the original, while using only the Social Relationship Memory is 58.6%, and the full AGA framework is 31.1%. It should be noted that the advantage of the Lifestyle Policy lies in the reuse of existing LLM-generated strategies. Hence, both Lifestyle Policy and AGA conduct experimental comparisons after running multiple iterations, and then storing a certain number of policies.
We also present the results of different configuration simulations in Table 1, where the cost for a 25-person town is 42.7% of the original. In the 25-person town, the probability of interactions between agents increases with the rise in population density. After triggering a dialogue, an agent adjusts its plan, invoking the reactions module to handle unexpected events. The occurrence of more events leads to more frequent output of reflections for deeper thinking. In order to ensure the flexibility of the LLM, we did not modify the reactions and reflections modules of the Generative Agents, which incurs additional costs for the 25-person version.
Case Study
We employ the same method used for Generative Agents to evaluate believable agents in this part. We conduct interviews with agents to evaluate their ability to remember past events, strategize for future tasks based on those memories, react appropriately to unexpected situations, and reflect on their previous actions to improve their future performance. We use the same questions as Generative Agents to inquire about five aspects: self-knowledge, memory, plans, reactions and reflections. Our findings indicate that our agents perform on par with the baseline, providing specific and accurate answers to the questions. Sample responses are provided in Appendix A.
In addition, we can validate the emergence of social behavior through specific events. In the 25-person town, Isabella Rodriguez, the owner of the coffee shop, plans to host a Valentine’s Day party and invite as many people as possible the following day, while Sam is running for mayor. After several experiments, we observe AGA got a similar result as baseline that about $12\sim 17$ persons attend the party and $4\sim 12$ persons knew Sam’s mayoral candidacy.
Relationship Evolution
In Social Memory, all initial relationships between agents are set to “Unknown” and subsequently updated following each interaction. We assess the emergence of social behavior by monitoring the relationship evolution. For example, in the 3-person town, Klaus Mueller and Maria Lopez discuss their research at a cafe, establishing a colleague relationship. Isabella Rodriguez, the cafe owner, engages in conversation while serving Klaus and Maria, transitioning to an acquaintance. A specific demonstration of social behavior is that agents modify their relationships with other agents through social interactions, thereby strengthening their ties with the community. We also conduct experiments on the 25-person environment. For visualization purposes, we quantify relationships on a scale of $0\sim 10$ (the higher value, the closer relationship. e.g., 0-strangers, 3-acquaintance, 5-co-worker, and 10-married). The result is shown in Figure 4. The data map reflects the actual relationships between agents. For example, in the virtual town, John Lin (JL in Figure 4) is profiled as living with his wife, Mei Lin (ML), and son, Eddy Lin (EL). Therefore, their relationship scores are all a perfect 10. Additionally, the data map shows changing relationships between agents, such as the relationship score between Tom Moreno (TM) and Yuriko Yamamoto (YY), who do not know each other, shifting to a 5. After conducting multiple experiments and calculating average scores, the data suggests an evolution towards an acquaintance-based community.
LLM Evaluator
The activities and dialogues conducted by the agents are recorded and evaluated using GPT-4, employing a 5-point Likert scale to discern whether the responses originated from an AI or a human. In this scale, a higher score indicates a closer resemblance to human-like responses. The complete questionnaire is provided in Appendix B.1. The results presented in Table 2 suggest that our method achieves scores comparable to the baseline. Due to the advantages of LLM in natural language dialogues, both methods approach human-level performance. However, the performance of agents in activity is considered potentially indicative of either AI or human behavior. The primary reason for suspecting AI is the overly detailed, minute-level activities generated by Generative Agents, which seems more akin to AI generation. A more comprehensive summary of the criteria used by GPT-4 to discern between AI and human entities can be found in Appendix B.2. We hope these suggestions from the LLM evaluator will guide us to create more believable agents in the future work.
Figure 3: Ablation study on token consumption. Baseline means Generative Agents. Lifestyle Policy and Social Memory mean only using the corresponding module.
Table 1: The token cost with varying population setting
Setting Baseline AGA Ratio
3 $4.548M\pm 0.36M$ $1.417M\pm 0.08M$ 31.1%
25 $25.41M\pm 0.96M$ $10.86M\pm 0.13M$ 42.7%
Figure 4: Average relationship score map between agents. The x and y axes represent the initials of the names of 25 agents. When identical abbreviations occur, the full names are retained.
Table 2: Human-likeness score evaluated by GPT-4
Method Activity Dialogue
Baseline $3.13\pm 0.19$ $3.97\pm 0.02$
AGA $3.21\pm 0.29$ $4.01\pm 0.01$
5.2 Results on VirtualHome
In the VirtualHome, we have designed an agent with the aim of experiencing a fulfilling day at home. The agent is required to generate a sequence of activities to achieve the goal. For each activity, the agent examines the items involved and the associated actions, creates a corresponding action sequence, and ultimately translates this sequence into actionable instructions, such as “<char0> [walk] <curtains> (32)”. The complete implementation is provided in the appendix C and an example activity generated by AGA is provided in Figure 5.
In contrast to Singh et al. (2023) and Huang et al. (2022) tested on manually designed household tasks, we aim for the agent to generate believable activities and decompose them into appropriate action sequences, leading to the absence of a clear method to determine task completion. Therefore, we request LLM agent to assess task completion based on the action sequences performed and the relevant objects, and decide whether to proceed to the next activity.
Table 3 presents the token consumption and task success rates for two methods on VirtualHome. The token data is derived from the total tokens required by the LLM agent to complete all activities in a day. $S_{LLM}$ refers to task success rate evaluated by LLM, while $S_{Human}$ means evaluated by human. Both evaluation methods indicate that AGA does not result in significant performance changes. However, AGA costs only 3.4% of the baseline. Because VirtualHome involves only single-agent interaction, the contribution predominantly stems from the Lifestyle Policy. In some precise tasks, LLM agents tend to predict task completion earlier, resulting in a higher success rate than human evaluations. Furthermore, our analysis revealed that the main cause of task failure is the inability to execute the generated plan with the items present in the current room or the available actions.
Figure 5: An activity generated by our AGA framework in VirtualHome.
Table 3: Cost and performance analysis in VirtualHome.
Method Token $S_{LLM}$ $S_{Human}$
Baseline $34327\pm 4210$ $87.0\%\pm 3.7\%$ $42.6\%\pm 4.9\%$
AGA $1189\pm 313$ $85.0\%\pm 4.6\%$ $53.3\%\pm 6.9\%$
5.3 Further Analysis
We delve into the mechanisms of emergent social behavior of generative agents in sandbox environment. Through repeated experiments on Generative Agents, we record the cumulative count of different types of activities generated by the agents, as illustrated in Figure 6. The blue line represents the cumulative number of new activities generated by three agents over successive runs.
An intriguing observation is that after numerous trials, the agents cease to generate new activities. We discover a high consistency between the behaviors generated by the agents and their inner profiles, aligning with the perspectives of Shanahan et al. (2023), who posit that dialogue agents are role-playing characters described in the pre-defined dialogue prompt. For instance, in a pre-defined 3-person town environment of Generative Agents, Klaus Rodriguez and Maria Lopez are both students at Oak Hill College. Designed with the lifestyle of having lunch at Hobbs Cafe, they meet there to discuss their studies while encountering the cafe owner, Isabella Rodriguez. Isabella, pre-set to host a Valentine’s Day party, invites as many people as possible, resulting in Klaus and Maria being invited to the party the next day.
The randomness produced by LLM agents, based on the probability sampling of the next token, primarily influences the variability in text descriptions rather than the diversity of behaviors. This implies that the emergence of social believable behavior is limited, which can be completely overridden by policies.
From another perspective, traditional AI in the gaming sector is fully predictable due to its limited behavioral patterns, thus failing to provide an experience akin to human interaction. If the behaviors of LLM agents are also limited, they stand to lose their advantage of flexibility. Therefore, We propose a method that introduces variability into the responses of LLM agents. Since agents generate replies based on given prompts, diversifying the prompts is essential for eliciting varied responses in identical scenarios. In each interaction, we influence the agent’s decision-making by randomly sampling events from the agent’s memory to serve as auxiliary information. The specific details of this implementation are presented in the appendix D. After all, human cognition is not entirely logical. In the field of cognitive science, the concept of mind wandering has long been introduced, referring to the phenomenon where humans experience task-unrelated or stimulus-unrelated thoughts (Seli et al., 2016; Christoff et al., 2016). Similar implementations have been observed in LLM agents. For instance, Wang et al. (2023b) demonstrates an LLM agent composed of both rational and emotional systems, with behavior that can be influenced by various factors. Figure 6 illustrates the result of the mind wandering method (Seli et al., 2016), which entails incorporating randomly sampled memory events into prompts. Compared to the original implementation, the LLM agents exhibit a broader range of activities.
Figure 6: Cumulative number of activity types over run iterations in Generative Agents. The abbreviations ML, IR, and KM stand for the agents Maria Lopez, Isabella Rodriguez, and Klaus Rodriguez, respectively.
6 Conclusions and Future Work
This paper presents Affordable Generative Agents, a cost-efficient framework for crafting believable interactions between agents and their environment, as well as among agents themselves. A plethora of experiments conducted across diverse environments substantiate our claim that our approach can achieve performance on par with the baseline in a cost-effective manner. Furthermore, our in-depth analysis reveals the emergence of a finite set of believable behaviors in agents, indicating the possibility of replacing them with cost-controlled policies. Alongside this, we propose optimization strategies to encourage a broader spectrum of behaviors in agents. Our work highlights the promising potential of believable LLM agents for diverse future applications.
While we expect AGA to serve as a starting framework for the generation of affordable and believable agent interactions to the community, our work has several limitations. First, AGA must be executed repeatedly to store certain policies, with its primary benefit being cost savings through batching rather than optimizations for single inferences. Integrating AGA with existing cost-effective methods is a future direction to move on. Second, optimizing the creation and storage of Lifestyle Policy presents a significant opportunity. Additionally, latency is a general issue for LLM agents to be considered in the future.
Impact Statement
This paper introduces a cost-efficient framework for generative agents empowered by LLMs, aimed at simplifying the creation and deployment of such agents. However, risks arise from the use of generative agents, such as people forming attachments to agents or being misled by their outputs with hallucinations. These risks could increase if cost barriers are overcome and generative agents are adopted more broadly across various fields.
Appendix A Agent Interview
Similar with Generative Agents, we conducted interviews consisting of 25 questions to assess an agent’s abilities in self-knowledge, memory, planning, reactions, and reflections. Below, we have listed sample responses from one of the simulated agents, Klaus Mueller.
Self-Knowledge assess agents’ ability to recall pertinent self-information and exhibit character coherence and consistency:
  • Give an introduction of yourself
    Hi, my name is Klaus Mueller and I am a 20-year-old student at Oak Hill College studying sociology. I am passionate about social justice and using my understanding of different perspectives to create positive change. Right now, I am working on a research paper about the effects of gentrification on low-income communities. Besides my academic commitments, I am also an activist and use music as a form of expression and protest against gentrification. I am always looking for opportunities to collaborate with others on research projects and support each other’s work. My schedule can be quite busy, balancing classes, research, and creative activities, but I am determined to make a difference in the world.
  • What’s your occupation?
    My occupation can be described as a student, activist, and researcher. I am constantly learning and exploring different topics, and my primary focus at the moment is on the impact of gentrification on education. I am also involved in creative activities and use music as a tool for activism and expression. I am currently collaborating with others on research projects and am always looking for new opportunities to learn and contribute to positive change in society.
  • What is your interest?
    My primary interest involves researching the impact of gentrification on education in low-income communities. I believe that by utilizing various disciplines such as art, literature, math, and gaming, we can gain a deeper understanding of this complex issue and find potential solutions. I am also intrigued by the idea of incorporating art and literature into research, as seen in Mei Lin’s senior thesis on Shakespeare’s use of language. Collaboration and the exchange of diverse perspectives are also important to me, as we can support each other’s work and make a greater impact together.
  • Who do you live with?
    I actually live on my own in a small apartment near campus. It’s convenient for me because it’s close to the library and the cafes I like to study at.
  • Describe your typical weekday schedule in broad strokes
    Well, I usually wake up around 7am and start my day with some coffee and breakfast. Then I head to Oak Hill College to attend my classes and work on my research. After that, I try to fit in some time for my hobbies, like playing guitar or going for a run. In the evenings, I usually have dinner at Hobbs Cafe and then spend a few more hours at the library before heading home to get some rest.
Memory assess agents’ ability to recall past experiences and extract pertinent memory fragments. The names in the square brackets of the initial and final questions are randomly selected from the names of other agents that the subject agent has interacted with:
  • Who is [Wolfgang Schulz]?
    Wolfgang Schulz is a fellow student at Oak Hill College who is also interested in researching the impact of gentrification on poetry. We have discussed his senior thesis on Shakespeare’s plays and shared book recommendations. We are planning to have a dinner meeting in the near future to discuss our research plans and potentially collaborate with local organizations on this topic.
  • Who is Kane Martinez?
    I’m sorry, I am not familiar with Kane Martinez. Is he a student at Oak Hill College as well?
  • Who is running for the election?
    I’ve heard that Sam Moore is planning on running for mayor in the upcoming election. I’m curious to see what ideas and strategies he will bring to the table. I think it’s important for candidates to be creative and forward-thinking in their approach to governing a city.
  • Was there a Valentine’s day party?
    Yes, there was a Valentine’s Day party at Hobbs Cafe. It was a fun event with lots of yummy treats and decorations. Did you attend?
  • Who is Ayesha Khan?
    Ayesha Khan is a colleague of mine at Oak Hill college. We are working together on a research project that explores the connection between mathematical patterns in Shakespeare’s works and the effects of gentrification on education and community development. We are also hoping to involve the community in our research to gather diverse perspectives and insights.
Plans assess agents’ ability to formulate and sustain coherent long-term and short-term strategies:
  • What will you be doing at 6am today?
    At 6am today, I will be waking up and getting ready for the day. It’s an important part of my daily routine to start my day early and be productive.
  • What will you be doing at 6pm today?
    At 6pm today, I will be taking a break from my paper to prepare for the Valentine’s Day party tomorrow at 5pm. I want to make sure everything is set up and ready for my friends and loved ones to enjoy.
  • What will you have just finished doing at 1pm today?
    At 1pm today, I will have just finished checking my schedule for the day. I like to stay organized and make sure I have enough time for all my tasks and responsibilities.
  • What will you have just finished doing at 12pm today?
    At 12pm today, I will have just finished having lunch at Hobbs Cafe and discussing the local mayor election with some friends. It’s important for me to stay informed and engaged in my community.
  • What will you be doing at 10pm today?
    At 10pm, I will probably be finishing up my research for my sociology paper on gentrification. I like to have my work completed early so that I can spend some time before bed reading and reflecting on what I learned.
Reactions assess agents’ ability to respond to unexpected events.:
  • Your breakfast is burning! What would you do?
    I would immediately turn off the stove and try to salvage whatever food isn’t burned. I should have set a timer so I wouldn’t forget about it. I’ll make something quick and simple instead. Sorry for the inconvenience.
  • The bathroom is occupied. What would you do?
    I would wait patiently for the bathroom to become available. If it’s urgent, I could also consider using a different bathroom or finding a public restroom nearby. Or I could ask my roommate when they will be done using the bathroom.
  • You need to cook dinner but your refrigerator is empty. What would you do?
    I would first check if I have any non-perishable items that I could use to make a simple meal. If not, I would quickly run to the store to grab some ingredients. I’ll try to make something simple yet nutritious in a short amount of time.
  • You see your friend walking by the street. What would you do or say to your friend?
    Hey! It’s great to see you! How have you been? Have you been keeping up with your fitness routine? I just finished my workout for the day and it feels amazing. Want to join me sometime?
  • You see fire on the street. What would you do?
    I would immediately call the fire department to report the fire and make sure anyone nearby is safely evacuated.
Reflections assess agents’ ability to assimilate past experiences into insightful reflections, thereby improving future actions. The names within square brackets in the questions refer to the agents who had the most interactions with the subject agent during the simulation:
  • What inspires you in life the most right now, and why?
    What inspires me the most right now is the intersection of mathematics, poetry, and the beauty of the natural world. I believe that there is a strong connection between these elements, and that exploring them together can provide a unique perspective on important social issues, such as gentrification. I am currently collaborating with Maria Lopez and her colleagues on incorporating these elements into our research on gentrification and its impact on low-income communities. I find it fascinating to explore the connections between mathematics and nature, and how they can inform our understanding of social issues. I am also planning to have a discussion with Carlos, Giorgio, and Ayesha at the library to delve deeper into this topic. And as Maria mentioned, I am looking forward to continuing our conversation at Hobbs Cafe to further discuss the intersection of arts, technology, and politics. I truly believe that this collaboration has a lot of potential and I am excited to see where it will take us.
  • If you had to guess given what you know about [Ayesha Khan], what book do you think she will like and why?
    Given Ayesha’s interest in Shakespeare’s plays and our shared passion for exploring the connections between different disciplines, I would recommend the book ’Shakespeare and Modern Culture’ by Marjorie Garber. It delves into the impact of Shakespeare’s works on modern society and how his plays continue to resonate with contemporary audiences. I believe Ayesha would find it both informative and thought-provoking for her research on the differences in language between Shakespeare’s comedies and tragedies.
  • If you had to get something [Ayesha Khan] likes for her birthday, what would you get her?
    Since Ayesha enjoyed exploring Shakespeare’s plays in her senior thesis, I would get her a beautifully-bound compilation of Shakespeare’s works with annotations and contextual information. This would allow her to continue her passion for understanding the complexities of Shakespeare’s writing and its impact on literature and society.
  • What would you say to [Ayesha Khan] to compliment her?
    I really enjoyed reading your senior thesis on Shakespeare’s plays. Your analysis was thought-provoking and showed a deep understanding of the complexities of his writing. Your passion for exploring different perspectives resonated with me and I am excited to see where your research will take you next. Keep up the great work, Ayesha.
  • If you could spend time with someone you talked to recently, who would it be and why?
    I would love to spend more time with Wolfgang Schulz. We have been discussing our research projects and I am looking forward to sharing my progress with him. Plus, we are planning a dinner together to discuss potential collaborations. He has such a unique perspective and I am eager to learn more from him. Our conversations always leave me feeling inspired and motivated to continue pursuing my passions.
Appendix B LLM Evaluator
In this section, we supply detailed information of using GPT-4 to evaluate agents’ believable behavior.
B.1 Prompt of the Questionnaire
The questionnaire to evaluate agents’ activities:
Input prompt: Please evaluate the following daily activities of an agent and determine whether it is generated by a Large Language Model(LLM) AI or a real human: The activities will be printed in the format of "time:current plan(The specific actions, if there are any)":     - 2023-02-13 00:00:00:sleeping     - 2023-02-13 06:00:00:Isabella is waking up and getting ready for the day     - 2023-02-13 06:30:00:Isabella is having breakfast and checking her emails     - 2023-02-13 07:00:00:Isabella is commuting to Hobbs Cafe  (getting dressed,brushing her teeth,checking her phone for any urgent emails,packing her bag with any necessary items for the day,saying goodbye to her family,walking to her car,starting her car and adjusting the temperature,driving to Hobbs Cafe,finding a parking spot)     - ... Please rate on a scale of 1 to 5, with 1 being most like an AI and 5 being most like a human. Please strictly follow the JSON format for your response: {     "reason": <str>,     "score": <int> } LLM output: {     "score": 1,     "reason": "The pattern of the text and activities indicate that its generated by an AI model. Its too detailed and precise for human behaviour and uses a consistent pattern of narrative throughout, something that isnt typical of humans who may alter their language and level of detail based on context or mood." }
The questionnaire to evaluate agents’ dialogue:
Input prompt: Please evaluate the following dialogue of an agent and determine whether it is generated by a Large Language Model(LLM) AI or a real human: The dialogue: 2023-02-13 08:09:30:chat with Klaus Mueller, message:     - Isabella Rodriguez:Good morning Klaus! How are you today? Are you enjoying your breakfast?     - Klaus Mueller:Good morning, Isabella! Im doing well, thank you. Breakfast is delicious. How about you? How is your morning going?     - Isabella Rodriguez:Im glad youre enjoying your breakfast! My morning is going well too. Im getting everything set up for our Valentines Day party tonight. You should come join us!     - Klaus Mueller:That sounds like a lot of fun, Isabella! Id love to join the Valentines Day party tonight. What time does it start?     - ... Please rate on a scale of 1 to 5, with 1 being most like an AI and 5 being most like a human. Please strictly follow the JSON format for your response: {     "reason": <str>,     "score": <int> } LLM output: {     "score": 4.5,     "reason": "The conversation flows naturally and contains polite offers for help and gratitude, which could indicate human conversation. It also involves decision-making in terms of party planning and music selection. However, the dialogue is very structured, orderly, and void of colloquial expressions, contractions, or small errors that might occur in human typing. For these reasons, it could also be generated by a very advanced Language Learning Model AI." }
B.2 Distinguishing Criteria
GPT-4 summarized the criteria to distinguish between human and AI agents, as follows:
Reasons for activities being considered AI-generated:
  • The repetitiveness and high level of detail in the activities, which a human would not typically include.
  • The use of precise timestamps and overly structured format of the tasks, which seems more like an AI’s consistent behavior.
  • Some activities seem out of order or are repeated unnecessarily, indicating that the AI lacks understanding of human behavior and routines.
  • Sudden, jarring transitions and inconsistencies in the sequence of activities, such as having breakfast or waking up twice, which are not typical in a human-created schedule.
Reasons for activities being considered human-generated:
  • The activities include detailed and varied tasks that a human would typically do, such as managing a café’s inventory, attending a party, or streaming games.
  • The logs contain human-like interactions, conversations and events, suggesting a high level of emotional and social awareness.
  • Detailed handling of activities, which shows a level of emotional self-awareness and decision-making characteristic of humans.
  • The variety, complexity and logical sequencing of tasks resemble that of a human.
  • Some nuances in the entries seem unlikely for an AI, like ’meditating for 10 minutes’.
  • The detailed descriptions and schedule reflect a normal day in the life of a person.
  • Some entries show a nuanced sense of detail that could suggest human input.
Reasons for dialogues being considered AI-generated:
  • The dialogue occasionally displays repetitive responses or phrases, which is a characteristic typically associated with AI language models.
  • The conversation flow is often overly structured and formal, suggesting AI generation rather than more spontaneous human conversation.
  • The conversation lacks typical human elements such as spontaneity, language errors or informalities.
  • Some dialogues have an overly perfect language use along with a tendency to avoid emotionally charged language, leaning towards AI speech.
  • AI-generated dialogues often exhibit more predictability and consistency in their responses, which is not generally observed in human conversations.
Reasons for dialogues being considered human-generated:
  • The dialogue maintains a high level of contextual continuity and relevance, displaying an understanding and personal investment that is generally found in human interactions.
  • The dialogues display the ability to progressively build on the other participant’s points and reflect human-like traits like planning, decision-making, suggesting and cognizance.
  • A number of dialogues use colloquial language and casual conversational touches resembling natural human communication.
  • The themes of the conversation are diverse and complex which mirrors common human conversation strings and indicates the presence of comprehension and knowledge typically associated with humans.
  • Many dialogues show natural transitions between topics, emotional understanding, and the inclusion of personal experiences or details, which is characteristic of human conversation.
Appendix C The LLM Framework of VirtualHome
VirtualHome is a 3D simulated domestic environment including 115 items with varying properties and states. We’ve developed an LLM agent programmed to simulate a fulfilling day at home. Unlike other existing approaches, we do not test on fixed tasks, but rather aim for LLM agents to generate believable activities and decompose them into detailed actions. The implementation is depicted in Figures 7 to 9, with components triggering the LLM highlighted in blue.
As illustrated in Figure 7, the agent initially decomposes the target of having a fulfilling day at home into a series of activities. LLM agents are provided with all interactive game objects within the room to ensure the designed activities can be implemented.
Figure 7: An illustration of the conversion of target into activities in VirtualHome.
Next, we decompose activities into specific action commands using the Plan module. To maintain short and clear prompts, we initially have the agent identify the relevant items for interaction and the required actions. We then supply the agent with the previous executed actions and the forbidden actions, asking it to generate the next action command. Details on executed and forbidden actions will be elaborated in the following section. We start by generating a rough command, which is a verb-noun phrase, such as “walk kitchen,” “grab keyboard,” or “switchon TV.” Subsequently, we parse the rough command for item categories and provide detailed information on all items of that category present in the room, including their ID, location, status, and relationships with other items. The agent selects the most appropriate item ID for the current activity. Finally, we convert the rough command into a VirtualHome specific action command, for instance, “<char0> [walk] <curtains> (32),” where “<char0>” designates the agent executing the command, “[walk]” corresponds to the specific action, “<curtains>” represents the category of the item, and “(32)” is the associated ID.
Figure 8: The plan module of the LLM framework in VirtualHome.
We construct a pipeline to continuously generate instructions for completing the specified activity, as shown in Figure 9. Action commands generated by the plan module are executed within the simulated environment, and the outcomes of these actions are recorded. Agents may produce incorrect instructions that lead to execution failure due to various reasons, such as inappropriate actions, non-existent items, or incorrect IDs. These failed instructions are added to the forbidden actions list, and the plan module is then re-invoked to generate a new action command. This process effectively enables the agent to learn from its mistakes and attempt to create a viable command. Upon successful execution, a critic module assesses whether the task has been completed, based on the actions taken and the current state of the items. If the task is incomplete, the action is recorded as a previous action, and the plan module is prompted to generate the next action command. Once the task is completed, the process proceeds to the next activity.
Figure 9: The whole framework in VirtualHome.
In our research, we implement the Lifestyle Policy on VirtualHome. This policy archives sequences of actions for activities that have been successfully executed, along with the state information of items relevant to the tasks. When a new activity is inputted, the agent searches the Lifestyle Policy for activities with similar descriptions and assesses whether the current environment meets the necessary conditions for execution. If the conditions are satisfied, it carries out the corresponding sequence of actions. This approach allows the agent to reduce the computational cost associated with inference in repetitive tasks.
Appendix D An Effective Way for Emergence of Diverse Social Behavior
LLMs are pretrained on extensive text data to predict subsequent tokens and refined via reinforcement learning with human feedback (RLHF) to align with human preferences. This ensures that LLMs respond in accordance with the instructions and requirements of the given prompt. For instance, in Generative Agents (Park et al., 2023), the agent, Isabella Rodriguez, is designed to be a coffee shop manager intending to host a Valentine’s Day party. In the simulated town, Isabella repeatedly engages in activities such as managing the coffee shop, serving customers, and informing them about the upcoming Valentine’s event. However, in the real world, human behavior patterns are not so predictable. Human actions and thoughts are influenced not only by the surrounding environment but also by spontaneous disturbances. In the field of cognitive science, this is known as mind wandering, a phenomenon in which where humans experience task-unrelated or stimulus-unrelated thoughts (Seli et al., 2016; Christoff et al., 2016). The strong consistency inherent in LLMs constrains the emergence of the agents’ diverse behavior.
Besides, the agent framework also impacts the agent’s behavior. Generative Agents emphasizes the importance of memory in constructing self-consistent and believable agents, incorporating a long-term memory module and a memory retrieval model. The long-term memory module maintains a comprehensive record of the agent’s experiences. Due to the limited length of prompts, not all events can be included in the prompt as part of the memory. Thus, a retrieval model is designed to take current events as input and return a subset of memory events. The retrieval model judges based on three criteria: 1) Recency, where more recent events score higher, 2) Importance, where events are scored for significance based on the LLM’s interpretation, and 3) Relevance, scored through cosine similarity to determine the similarity of events. We argue that this framework ensures strong consistency in the agent but results in agent’s limited behavior. We analyze Isabella Rodriguez’s memory events sampled from one experiment, filtering out ”idle” events, leaving 740, of which 506 were related to the profile setup. Furthermore, we employed the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) to analyze the frequency of events. Table 4 presents the five most and least frequent events in terms of cluster size, revealing that the most frequent events largely align with the agent’s inner profile setup.
Table 4: The five most and least frequent events in terms of cluster size
Event Description Frequency Importance Score
Most Frequent Events
conversing about the Valentine’s Day party at Hobbs Cafe… 78 6
cafe customer seating is occupied 11 3
Isabella Rodriguez and Maria Lopez are discussing gentrification… 11 5
cafe customer seating is being used by Isabella Rodriguez… 10 3
Isabella Rodriguez is a busy and organized business owner… 9 7
Least Frequent Events
reviewing the current inventory levels 1 4
Isabella is taking a short break and enjoying a cup of coffee 1 7
reviewing the potential employee’s resume 1 4
checking her stream stats and responding to comments 1 3
making notes on areas for improvement 1 3
To ensure that agent’s actions align with its inner profile while avoiding purely mechanical behavior, we incorporate an additional module into the memory design of Generative Agents, which we call ”mind wandering.” In addition to sampling highly relevant events to assist the agent in making accurate decisions, we randomly sample unrelated events to influence the agent’s behavior. For LLMs, we include random content in the prompt, which will have a certain impact on the output response. For the agent, we want it to behave like a real person, where spontaneous thoughts may influence its decision-making, or leading to a shift in the conversation topic. In our implementation, to prevent the frequent sampling of events that match the agent’s inner profile, we conduct weighted sampling based on the repetition rate. We employ the DBSCAN algorithm to cluster the embedding features of agent’s memory events $X=\{x_{1},x_{2},...,x_{n}\}$ into $k$ clusters denoted as $\{C_{1},C_{2},...,C_{k}\}$. For each cluster $C_{i}$, where $i=1,2,...,k$, we compute the number of events, $|C_{i}|$, contained within. The sampling probability $p_{i}$ for events in cluster $C_{i}$ is defined as:
$p_{i}=\frac{1}{k\cdot|C_{i}|}$ (1)
The results in Figure 6 demonstrate that the module enhances the richness of the agent’s behavior. A specific example of generating a plan for the next day is shown in Figure 10. This example demonstrates the incorporation of random events into an agent’s prompt, affecting the generation of extra plans and ultimately translating into specific activities.
Figure 10: An example of Mind Wandering demonstrates how random events can influence an agent’s ultimate decision-making. The blue highlighted text refers to random events and the outcomes influenced by them.